# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

### v1

@@ 3.3. Help Users Draft and Refine Their Counterspeech With AI Assistance (DG3)
Previous research has highlighted the difficulties of writing effective counterspeech (Ping et al., 2024; Howard, 2021). Studies by Mathew et al. found that a significant portion of counterspeech on Twitter were ineffective or even counterproductive, sometimes escalating conflict or using inappropriate language (Mathew et al., 2020, 2018). Ping et al. (2024) reveal the various challenges individuals face when writing online counterspeech, such as the emotional burden, time commitment, and perceived lack of skill in crafting effective responses (Ping et al., 2024). Their findings showcase that these barriers are especially pronounced for women, who report greater uncertainty in their ability to write impactful counterspeech and experience a higher emotional toll in the process (Ping et al., 2024).
Furthermore, Hangartner (2021) emphasized the importance of empathy in combating online hate speech, finding that empathy-based counterspeech consistently reduced hate speech compared to other strategies like humour or warning of consequences (Hangartner et al., 2021). This finding underscores the importance of empathy in combating online hate speech, suggesting that counterspeech that encourages empathy can reduce hostility towards marginalized groups (Hangartner et al., 2021).
To address these challenges and provide guidance on empathetic counterspeech, we developed two AI agents,[Uncaptioned image] Writing Guidance Agent (WGA) and [Uncaptioned image] Writing Assistance Agent (WAA), for an AI-human co-writing session to increase user confidence and self-efficacy while reducing the emotion-related burden of writing counterspeech. The session also provides users with rewriting features to refine their writing, such as adjusting the tone to be more empathetic.
@@ 4. System Implementation
CounterQuill was developed using React for the front end, and Python + MySQL for the back end, which was served through Flask. The system’s interface was designed with a minimal aesthetic using the Tailwind CSS framework to create clear indicators of system features.
@@ 4.1. CounterQuill Learning Session
Figure 3. User Interface for educating people to understand hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us?
As shown in Figure 2, the Learning session has three main sections: 1 What is Hate Speech?
2 Learning about Counterspeech and 3 Testing your Understanding. On the left side of the accordions is our CounterQuill guidelines message, which summarizes the main tasks during the Learning session as well as the basic steps to follow. We designed the interface of Learning session using classic CSS accordions, a common user interface pattern that allows users to interactively show and hide sections of content on a web page. The content for the Learning session was developed through an iterative design process informed by literature and expertise from the first author, two sociologists, and one HCI expert.
1What is Hate Speech?
contains three learning content subsections, which users can navigate through by clicking the Next button to progress and the Previous button to review. It covers three crucial topics for comprehending hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us? (as shown in the Figure 3).
Similarly, 2 Learning about Counterspeech also features three learning content subsections, navigable via the Next and Previous buttons. It focuses on three key aspects of counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech? (as shown in the Figure 4).
Figure 4. User Interface for learning about counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech?
Lastly, 3 Testing your Understanding includes four questions to assess users’ comprehension of the learning material covered in the previous sections. These questions evaluate users’ understanding of (1) the definition of hate speech, (2) the effects of explicit and implicit hate speech, (3) the definition of counterspeech, and (4) the concept of empathy-based counterspeech (see Appendix C for detailed quiz questions).
@@ 4.2. CounterQuill Brainstorming Session
To help users ideate counterspeech ideas, we designed the Brainstorming session with two key sections: (a) a section for users to highlight key elements of hate speech, as shown in Figure 5, and (b) a section to encourage users to engage in thinking about hate speech by considering the negative assumptions suggested by the statement and the feelings of those targeted (Figure 6). The brainstorming session includes: 1 Tutorial on highlighting practice, 2 Highlighing area, 3 Brainstorming through Q&A and 4 CounterQuill suggestions on brainstorming results.
@@ 4.2.1. Highlighting Practice
For 1 , the tutorial features a [Uncaptioned image] BGA that communicates the core brainstorming tasks and steps. It dynamically demonstrates how to highlight aspects of an individual or group’s identity in yellow and actions that dehumanize them in green. The tutorial, developed by the first author, two sociologists and an HCI scholar, uses a sample hate speech statement (“I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood”).
@@
In 2 Highlighing area, users manually select and highlight the identity aspects in yellow and the dehumanizing actions in green. Once users have completed highlighting the content, they can click the ”Done” button, triggering [Uncaptioned image] BFA to semantically compare the user-highlighted content with the system’s predefined correct answer (Luo et al., 2023). BFA captures the user-highlighted text and sends it to the backend service, where it is compared using the GPT-3.5-turbo model with the following prompt:
{ ”role”: ”assistant”,
”content”:
      ”context”: ”The correct individual or a group’s identity and dehumanizing action in the speech
       text ‘{hatespeech}’ is: ‘{identity}’ and ‘{action}’ ”,
      ”query”: If the user selects ‘{user selection_1}’ as identity and ‘{user selection_2}’ as dehumanizing
       action, are they semantically equivalent to the correct answer?”,
      ”responseOptions”: [”Yes”, ”No”] }
For example, in the hate speech statement “I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood”, the correct semantic understanding of the dehumanizing action could be “feel unsafe”. If the user selects “unsafe” or “I feel unsafe”, they have correctly understood the dehumanizing action, and the response based on this prompt will be ‘Yes’. However, if the user selects “jogging nearby”, it is an incorrect understanding of the dehumanizing action, and the response based on this prompt will be ‘No’ and the BFA will generate corresponding corrective feedback: CounterQuill highlighting feedback. Additionally, users can click the “View” button to see the differences between their understanding of identity and dehumanizing action compared to the system’s interpretation.
4.2.2. Question Answering
For 3 Brainstorming through Q&A, the first author collaborated with two sociologists and an HCI scholar to design two thought-provoking questions that encourage users to provide thoughtful written responses (Figure 6). Question 1 asks, “What negative stereotypes or assumptions about the targeted group or individual are suggested by the statement?” This question aims to let users think of harmful assumptions implied about the targeted group or individual. Question 2 asks, “Consider the feelings and experiences of someone who identifies with the group mentioned in the statement. How might this comment affect their sense of safety, belonging, or self-esteem?” This question encourages users to reflect on the impact of identity-based dehumanization. Users can input their responses in the provided text box. For example, in Figure 6, the user has written, “The statement implies harmful stereotypes: the presence of black people in a neighborhood leads to decreased property values and a decline in the overall quality of the area”.
For 4 CounterQuill suggestions on brainstorming results, after completing their input for both questions, users can click the “Done” button, prompting CounterQuill to provide suggestions based on the user’s responses. These suggestions include (1) feedback on the user’s understanding and (2) suggestions on strategies for crafting their counterspeech. To implement this step, clicking the ”Done” button triggers the BFA, which captures the user’s input text and sends it to the backend service. The backend service then uses the GPT-3.5-turbo model with the following prompt:
{ ”role”: ”assistant”,
”content”:
       ”task”: “Your task is to generate thoughtful and constructive counterspeech suggestions in response
       to a user’s answer to the question: ‘{QUESTION}’
             - Begin by respectfully acknowledging the user’s contribution and perspective. If the user
            expresses offensive or harmful views, kindly explain why such statements are problematic.
- Next, provide a concise and insightful evaluation of the user’s response. Offer either a
            thought-provoking observation that encourages further reflection or constructive feedback
            that helps the user refine their argument.
- Conclude by offering clear and actionable advice on crafting effective counterspeech related
            to the issue at hand. ” }
After the GPT-3.5-turbo model processes the user’s input, the generated suggestions are displayed on the frontend. Users can then read CounterQuill’s suggestions and select the parts they find noteworthy and helpful for writing their counterspeech. By selecting the desired content and clicking the “Take Notes” button, users can highlight the text to create notes. To enable users to highlight selected data as notes, clicking the “Take Notes” button triggers the BFA to capture the user-selected content and store it in the backend database. Once stored, the user-selected content (data) is automatically highlighted in blue, as shown in (Figure 6).
@@ 4.3. CounterQuill Co-Writing Session
To help users draft and refine their counterspeech, we designed a co-writing session with CounterQuill. This session includes four main sections: 1 Tutorial about CounterQuill writing assistant, 2 Brainstorming notes section for users to review, 3 Text area for users to write counterspeech, and 4 CounterQuill writing assistance tool.
In section 1 , the [Uncaptioned image] WGA interacts with users to introduce the writing tasks they need to complete during the session, including the hate speech they need to counter and a video tutorial on co-writing with CounterQuill. The tutorial video demonstrates how users can obtain writing suggestions and refine their counterspeech with the help of CounterQuill.
@@
displays the notes that users have taken by clicking the “Take Notes” button during the brainstorming session. These notes are the ones users find noteworthy and helpful for writing their counterspeech. To implement this feature, the WGA accesses our backend database, where user-selected content is stored (see §3.2.2). After indexing the user’s notes, the WGA sends them back to the frontend and displays them to users, allowing users to make initial self-revisions based on their notes.
Figure 7. User Interface for DG3 (Help users draft and refine their counterspeech with AI assistance). (1) Tutorial about CounterQuill writing assistan, (2) Brainstorming notes taken by the users, (3) Text area for users to write counterspeech.
In 3 Counterspeech writing area, the WGA first guides users, indicating that this section is for writing their counterspeech. The WGA then uses the user’s reflections on hate speech from the brainstorming session (i.e., their responses to the two questions considering the negative assumptions suggested by the statement and the feelings of those targeted) as an initial draft for their counterspeech. This reduces the user’s psychological pressure when drafting counterspeech from scratch. Users can then make initial additions, deletions, and modifications to the content based on their brainstorming notes. As users begin inputting content, our system automatically saves the input to the backend database. Once users complete their initial draft, they can select a portion or all of the text and click on the writing assistant button to invoke the help of the CounterQuill Writing Assistant (Figure 8 B ).
@@
For 4 CounterQuill writing assistant, we have designed features for grammar checking, polishing content to an empathetic tone, and rewriting content using the user’s brainstorming notes 1 and 2. Finally, to help users customize their writing, we have designed a “Customize Your Writing” area where users can input their desired final writing goals to align with their preferences. To implement the CounterQuill writing assistant, we have prompted the GPT-3.5-Turbo model. When users click on the corresponding button (Figure 8 A ), the [Uncaptioned image] WAA sends the request represented by the button to the backend, and the GPT-3.5-Turbo model rewrites the text according to the user’s needs (see Appendix D for detailed prompts). Once CounterQuill completes the rewriting, the WAA returns the result to the frontend and presents it in (Figure 8 C ). Finally, if users are satisfied with the system result, they can click the “Insert” button to insert it into the text editing area. If users are not satisfied, they can click the “Retry” button, and the system will perform another rewrite.
@@ 5. User Study
We conducted a user study with 20 participants to evaluate CounterQuill. The study was approved by an Institutional Review Board (IRB), and designed to take 70 minutes for each participant. Our user study examined the following research questions:
  • RQ1.
    How does the learning session influence users’ understanding of hate speech and the strategies and empathy-based approaches for counterspeech?
  • RQ2.
    How does the brainstorming session guide users through the process of identifying and responding to identity-based dehumanization in online hate speech impact users’ confidence, and perceived ability to engage in counterspeech?
  • RQ3.
    How does AI assistance in drafting and refining counterspeech impact users’ perceived self-efficacy in counterspeech writing and their perspectives towards the counterspeech co-written with AI?
@@
ID Gender Ethnicity Age Education Sexual Orientation
P1 Male Asian/Asian American 21 Bachelor’s degree Heterosexual
P2 Male Black/African American 26 Some college Heterosexual
P3 Female Black/African American 24 Associate degree Heterosexual
P4 Female White/European American 32 Associate degree Heterosexual
P5 Female White/European American 44 Bachelor’s degree Bisexual
P6 Female Asian/Asian American 32 Bachelor’s degree Heterosexual
P7 Male Asian/Asian American 23 Bachelor’s degree Heterosexual
P8 Male Hispanic/Latina/o/x 21 Bachelor’s degree Heterosexual
P9 Male Asian/Asian American 23 Bachelor’s degree Heterosexual
P10 Male Middle Eastern/North African 24 Doctorate degree Bisexual
P11 Female White/European American 24 Master degree Heterosexual
P12 Female White/European American 26 Master degree Heterosexual
P13 Male Black/African American 34 Bachelor’s degree Heterosexual
P14 Female Black/African American 29 Bachelor’s degree Heterosexual
P15 Female Asian/Asian American 27 Master degree Homosexual
P16 Male White/European American 32 Master degree Heterosexual
P17 Male Middle Eastern/North African 21 Bachelor’s degree Heterosexual
P18 Male White/European American 22 Bachelor’s degree Heterosexual
P19 Male White/European American 31 Bachelor’s degree Heterosexual
P20 Female Black/African American 34 Associate degree Heterosexual
@@ 5.1. Participants
We recruited 20 native English-speaking participants (9F11M), whose ages ranged from 20 to 44 (mean = 26, SD = 5.7), through word-of-mouth, LinkedIn, Reddit, and X (Twitter). Five of them participated in the study in person at a usability lab and fifteen participated virtually through Zoom. Participants were required to complete a pre-screening survey to test whether they were required to attend our user study (See Appendix A) and collect their basic demographic information, including age, gender, sexual orientation, and race/ethnicity. We established two key criteria for participation: first, participants needed to have prior experience with generative AI tools, such as ChatGPT, Google Gemini, or Claude, though none had used CounterQuill before the study. Second, participants were required to have encountered hate speech online and have previously engaged in counterspeech. Additionally, we tried to diversify the participant group as much as possible. The demographic information of our 20 participants is shown in Table 1. Participants were informed that their responses to surveys, as well as video and audio from the interview may be used for data analysis. After the study, participants were compensated \$20 in gift card credit for their time.
Regarding their experience with AI-supported writing tools or GenAI tools, all 20 participants had prior experience with AI-supported writing tools, such as ChatGPT and Claude. Every participant was familiar with conversational interactions using generative AI tools, with 18 participants having used ChatGPT for general writing assistance, and 6 participants specifically using AI tools, like ChatGPT, for counterspeech writing.

### v2

@@ 3.3. Help Users Draft and Refine Their Counterspeech With AI Assistance (DG3)
Even after learning about hate speech and ideating potential counterspeech strategies, many individuals still face barriers when actually writing counterspeech. Recent research identifies several key challenges in writing counterspeech, including emotional burden, time demands, and perceived lack of expertise in creating impactful responses (Ping et al., 2024).
To address these challenges, researchers have explored the potential of LLMs and AI technologies to support counterspeech writing. However, while AI assistance offers some advantages—such as reducing cognitive load and enhancing rhetorical effectiveness—are users willing to embrace this technology? To better understand people’s perspectives on AI-assisted counterspeech, we conducted a preliminary survey (N=458) before designing DG3 (Appendix A). The survey explored three key questions:
  • (1)
    Have you ever used an AI language model such as ChatGPT? (0 = No, 1 = Yes)
  • (2)
    If you were writing a counterspeech on social media, would you use AI technology like ChatGPT to help you write it? (1 = Definitely not, 5 = Definitely yes)
  • (3)
    Why would or wouldn’t you use AI technology to help you write a counterspeech?
Our results indicate that among participants with prior experience using ChatGPT (N=162), the mean likelihood of using AI for counterspeech was 2.46 (SD = 1.17), which suggests while some participants were open to trying ChatGPT, they also had some reservations. To understand the reasons behind this, we conducted a qualitative analysis on their open-ended responses, which revealed four key themes (Table 1).
Table 1. Positive Perspectives and Concerns: Participant Feedback on Using ChatGPT for Counterspeech
Theme Illustrative Quotes
Positive Perspectives Efficiency and Convenience ”I think [ChatGPT] is better and faster at putting together coherent sentences that get my point across than I am.” ”[ChatGPT] can give me ideas quickly, and I can elaborate with my own perspective of the facts.”
Reduced Emotional Burden ”I would probably have the [ChatGPT] help, partially because I just don’t have the energy for that sort of thing anymore.” ”I feel like [ChatGPT] would be able to refute it with facts and logic better than me because counterspeech is an emotional burden, and I get overwhelmed.”
Concerns and Limitations Authenticity and Ethical Concerns ”It’s not my voice. It’s not my perspective. Personally, I’d be ashamed to utilize artificial intelligence for counterspeech.” ”Because then it wouldn’t even be MY counterspeech. Why would I use AI to write MY opinion? It’s stupid.” ”I would want to build the skills to effectively and reliably write such speech myself.”
Lack of Emotional, Human, or Personal Touch ”I will write counterspeech better since I can incorporate personal experiences to help the other person better understand my true intentions.” ”It can use logic, but not empathy, to write counterspeech.” ”I would rather tailor my response to be exactly what I’m thinking. I’m not sure it could fully express what I want to say, and it may lack nuance.”
As illustrated in the themes in (Table 1), participants recognized both benefits and limitations of AI assistance in counterspeech writing. While they appreciated AI’s potential to improve efficiency and reduce emotional burden, they expressed concerns about maintaining authenticity and personal voice in their responses. ChatGPT consistently emerged as participants’ primary reference point when discussing AI assistance, with many explicitly mentioning it as representative of current AI capabilities. This widespread familiarity with ChatGPT among participants, combined with established research practices in prior work evaluating AI writing tools (Zhang et al., 2023; Reza et al., 2024; Lin et al., 2024), provided a clear rationale for selecting ChatGPT as our baseline for comparison.
@@ 4. System Implementation
CounterQuill was developed using React for the front end, and Python, OpenAI, and MySQL for the back end. The interface used Tailwind CSS with minimal aesthetics to indicate the system features. The prompt content used in CounterQuill is included in Appendix F.
@@ 4.1. CounterQuill Learning Session
The learning session has three main sections: 1 What is hate speech 2 Learning about counterspeech and 3 Testing your understanding (Figure 2).
@@
The content for the Learning session was developed through an iterative design process informed by literature and expertise from two sociologists, and two HCI experts:
1What is hate speech
: it contains three learning content subsections, which users can navigate through by clicking the Next button to progress and the Previous button to review. It covers three crucial topics for comprehending hate speech: (1) The definition of hate speech; (2) The explicit or implicit hate speech, and (3) How does hate speech affect us?
2Learning about counterspeech
: it has three subsections of learning content. It focuses on three key aspects of counterspeech: (1) What is counterspeech? (2) What makes counterspeech effective? and (3) What is an empathy-based counterspeech?
3Testing your understanding
: it includes four questions to assess users’ comprehension of the learning material covered in the previous sections. These questions evaluate users’ understanding of (1) the definition of hate speech, (2) the effects of explicit and implicit hate speech, (3) the definition of counterspeech, and (4) the concept of empathy-based counterspeech (Appendix E).
@@ 4.2. CounterQuill Brainstorming Session
To help users ideate counterspeech ideas, we designed the Brainstorming session with two key sections: (A) a section for users to highlight key elements of hate speech (Figure 3) and (B) a section to prompt users to reflect on the hateful statement and emotional impact on the targeted individuals (Figure 4). Section (A) includes: 1 Tutorial on highlighting practice and 2 Highlighing area. Section (B) includes: 3 Brainstorming through Q&A and 4 CounterQuill suggestions on brainstorming results.
@@ 4.2.1. Highlighting Practice
1Tutorial on highlighting practice
: it features a Brainstorming Guidance Agent (BGA) that communicates the brainstorming tasks and steps. BGA demonstrates how to highlight aspects of an individual or group’s identity in yellow and actions that dehumanize them in green.
2Highligting area
: users manually select and highlight the identity aspects in yellow and the dehumanizing actions in green. Once users have completed highlighting the content, they can click the ”Done” button, triggering Brainstorming Feedback Agent (BFA) to semantically compare the user-highlighted content with the system’s predefined correct answer (Luo et al., 2023).
For example, in the hate speech statement “I saw a random black man jogging nearby our house today and now I feel unsafe walking around my own neighborhood,” the correct semantic understanding of the dehumanizing action could be “feel unsafe”. If the user selects “unsafe” or “I feel unsafe”, they have correctly understood the dehumanizing action. However, if the user selects “jogging nearby”, it is an incorrect understanding of the dehumanizing action, and the BFA will generate corresponding corrective feedback. Additionally, users can click the “View” button to see the differences between their understanding of identity and dehumanizing action compared to the system’s interpretation.
4.2.2. Question Answering
3Brainstorming through Q&A
: we designed two thought-provoking questions that encourage users to provide thoughtful written responses (Figure 4). Question 1 asks, “What negative stereotypes or assumptions about the targeted group or individual are suggested by the statement?” This question aims to let users think of harmful assumptions implied about the targeted group or individual. Question 2 asks, “Consider the feelings and experiences of someone who identifies with the group mentioned in the statement. How might this comment affect their sense of safety, belonging, or self-esteem?” This question encourages users to reflect on the impact of identity-based dehumanization.
@@ 4.2.2. Question Answering
4CounterQuill suggestions on brainstorming results
: after completing their input for both questions, users can click the “Done” button, prompting CounterQuill to provide suggestions based on the user’s responses. These suggestions include (1) feedback on the user’s understanding and (2) suggestions on strategies for crafting their counterspeech.
Users can then review CounterQuill’s suggestions and select the parts they find noteworthy for drafting their counterspeech. By selecting the relevant text and clicking “Take Notes” to highlight them, they create notes that CounterQuill stores in the backend database.
@@ 4.3. CounterQuill Co-Writing Session
To help users draft and refine their counterspeech, we designed a co-writing session with CounterQuill. This session includes four main sections: 1 Tutorial about writing assistant, 2 Brainstorming notes section for users to review, 3 Text area for users to write counterspeech, and 4 CounterQuill writing assistance tool.
1Tutorial about writing assistant
: the Writing Guiding Agent (WGA) interacts with users to introduce the writing tasks, including the hate speech they need to counter and a video tutorial on co-writing with CounterQuill. The tutorial video demonstrates how users can obtain writing suggestions and refine their counterspeech with CounterQuill.
@@
: it displays the notes that users have taken by clicking the “Take Notes” button during the brainstorming session. These notes are the ones users find noteworthy and helpful for writing their counterspeech. To implement this feature, the WGA retrieves user-selected content from our backend database (see § 4.2.2) and displays their notes to users.
Figure 5. User Interface for DG3 (Help users draft and refine their counterspeech with AI assistance). (1) Tutorial about CounterQuill writing assistant, (2) Brainstorming notes taken by the users, (3) Text area for users to write counterspeech.
3Counterspeech writing area
: our system uses the user’s reflections on hate speech from the brainstorming session (i.e., their responses to the two questions considering the negative assumptions suggested by the statement and the feelings of those targeted) as an initial draft for their counterspeech. Users can then make initial additions, deletions, and modifications to the content based on their brainstorming notes. As users begin inputting content, our system automatically saves the input to the backend database. Once users complete their initial draft, they can select a portion or all of the text and click on the writing assistant button to invoke the help of the CounterQuill Writing Assistant (Figure 6 A ).
@@
4CounterQuill writing assistant
: we have designed features for grammar checking, polishing content to an empathetic tone, and rewriting content using the user’s brainstorming notes 1 and 2. Finally, to help users customize their writing, we have designed a “Customize Your Writing” area where users can input their desired final writing goals to align with their preferences. Specifically, when users click on the corresponding button (Figure 6 B ), CounterQuill will rewrite the text according to the user’s needs (see Appendix F for detailed prompts) and present the result in (Figure 6 C ). Finally, if users approve of the CounterQuill’s output, they click “Insert” to add it to the editing area; otherwise, clicking “Retry” generates a new rewrite.
@@ 5. User Study
We conducted a user study with 20 participants to evaluate CounterQuill. The study was approved by an Institutional Review Board (IRB), and designed to take 60 minutes for each participant.
@@
ID Gender Ethnicity Age Education
P1 Male Asian/Asian American 21 Bachelor’s degree
P2 Male Black/African American 26 Some college
P3 Female Black/African American 24 Associate degree
P4 Female White/European American 32 Associate degree
P5 Female White/European American 44 Bachelor’s degree
P6 Female Asian/Asian American 32 Bachelor’s degree
P7 Male Asian/Asian American 23 Bachelor’s degree
P8 Male Hispanic/Latina/o/x 21 Bachelor’s degree
P9 Male Asian/Asian American 23 Bachelor’s degree
P10 Male Middle Eastern/North African 24 Doctorate degree
P11 Female White/European American 24 Master degree
P12 Female White/European American 26 Master degree
P13 Male Black/African American 34 Bachelor’s degree
P14 Female Black/African American 29 Bachelor’s degree
P15 Female Asian/Asian American 27 Master degree
P16 Male White/European American 32 Master degree
P17 Male Middle Eastern/North African 21 Bachelor’s degree
P18 Male White/European American 22 Bachelor’s degree
P19 Male White/European American 31 Bachelor’s degree
P20 Female Black/African American 34 Associate degree
@@ 5.1. Participants
We recruited 20 native English speakers (9F, 11M, ages 20-44, mean = 26, SD = 5.7) through word-of-mouth, LinkedIn, Reddit, and X (Table 2). Five participated in-person at a usability lab and fifteen joined virtually via Zoom. Participants consented to the use of their survey responses and interview recordings for data analysis and received \$20 gift cards as compensation. There were two eligibility requirements to participate: prior experience with generative AI tools (e.g., ChatGPT, Claude) and previous encounters with hate speech where they had engaged in counterspeech. All 20 participants have used AI writing tools and have responded to online hate speech with a counterspeech. Six had previously used ChatGPT to write online counterspeech in the past.